# Gold-medalist Performance in Solving Olympiad Geometry with AlphaGeometry2

### v2

@@ Appendix C Featured AlphaGeometry2 solutions
The problem asks about the relationship between $\angle KIL$ and $\angle XPY$. The former is the angle formed by a midpoint and the incenter, which usually does not go well together and cannot be computed by the angles of the main triangle $ABC$. Typically, a human contestant would rely on trigonometry, complex numbers or other computational methods to find the solution. For AlphaGeometry, its DDAR system only relies on simple angle chasing and ratio chasing, so this necessitates the need for some auxiliary point constructions. To this end, AlphaGeometry constructs $E$ as a point on the line BI such that $\angle AEB=90^{\circ}$, which elegantly ties these seemingly unrelated geometric elements together by creating pairs of similar triangles $ABE$ and $YBI$, ALEALE and $IPC$. These pairs of similar triangles create new pairs of equal angles and equal side length ratio. That being said, the point $E$ gives purposes to the midpoint $L$ of $AB$. To complete the proof, we need to prove $\angle AIK=\angle BYP$ and $\angle AIL=\angle CPX$. To this end, we need to prove that triangle $AKI$ is similar to triangle $BPY$ and triangle $ALI$ is similar to triangle $CPX$, which is done by side length ratio chasing, which is obtained from the pairs of similar triangles above. A full solution is published at https://storage.googleapis.com/deepmind-media/DeepMind.com/Blog/imo-2024-solutions/P4/index.html. This solution was obtained within 30 seconds at IMO 2024 and was given the full seven points by Joseph Myers, a two-time IMO gold medalist and Chair of the IMO 2024 Problem Selection Committee.
@@
Along with IMO 2024 P4, AlphaGeometry can solve many challenging problems with only 1 extra auxiliary point, some of which involves rather unconventional constructions. One such problem is IMO 2013 P3.
@@
IMO 2014 P3: Convex quadrilateral ABCDABCD has $\angle ABC=\angle CDA=90^{\circ}$. Point $H$ is the foot of the perpendicular from AA to $BD$. Points $S$ and $T$ lie on sides ABAB and $AD$, respectively, such that $H$ lies inside triangle $SCT$ and $\angle CHS-\angle CSB=90^{\circ},\angle THC-\angle DTC=90^{\circ}.$ Prove that line $BD$ is tangent to the circumcircle of triangle $TSH$.
@@
Our final example is G7 problem from the IMO 2009 Shortlist.
@@
To the best of our knowledge, this problem previously had only computational solutions, e.g. by using complex numbers, trigonometric computations or proof by contradiction via an inequality argument. Since AlphaGeometry does not have access to these computational and reasoning tools, as well as advanced Euclidean geometry knowledge, we originally expected that this problem cannot be solved by AlphaGeometry. Nevertheless, AlphaGeometry was able to produce an elegant solution with only angle and ratio chasing by constructing key auxiliary constructions. First, AlphaGeometry shows that $X$ and $Z$ are reflections of each other w.r.t. $BI$, and by symmetry it follows that $I$ is the circumcenter of $\triangle XYZ$. From this we can show that $AB=AC$, and by symmetry we have $\triangle ABC$ is an equilateral triangle. However, the main challenge with this problem is to use the condition $\triangle XYZ$ being an equilateral triangle, i.e. $XY=YZ$ and its cyclic variants. To this end, AlphaGeometry constructs a series of circumcenters of key triangles:
@@
Appendix D Additional evaluation on the hardest IMO shortlist problems
@@
Appendix E Towards generating full proofs with a language model
@@
  • Unverified: The premises of the step does not imply the conclusion of the step under the deduction rule that it uses.
@@
For evaluation we query our language models with the 2000-2024 IMO problems and 32 samples at temperature 1.0. The models are queried without any end of sentence tokens such that they generate full proofs. Then we compute how many valid proof steps the models produce on average across samples and all problems. It turns out our models do not make many syntax mistakes (see Figure 16), the majority of generated steps are either valid (either fully verified or correct but unverified). One surprising find is that both small and larger models perform similarly. These results support ideas that large language models can be self sufficient without depending on external tools, but until inference speed is improved and hallucinations are completely resolved, the tools will stay essential for math applications.

### v3

@@ Appendix D Featured AlphaGeometry2 solutions
The problem asks about the relationship between $\angle KIL$ and $\angle XPY$. The former is the angle formed by a midpoint and the incenter, which usually does not go well together and cannot be computed by the angles of the main triangle $ABC$. Typically, a human contestant would rely on trigonometry, complex numbers or other computational methods to find the solution. For AlphaGeometry, its DDAR system only relies on simple angle chasing and ratio chasing, so this necessitates the need for some auxiliary point constructions. To this end, AlphaGeometry constructs $E$ as a point on the line BI such that $\angle AEB=90^{\circ}$, which elegantly ties these seemingly unrelated geometric elements together by creating pairs of similar triangles $ABE$ and $YBI$, ALEALE and $IPC$. These pairs of similar triangles create new pairs of equal angles and equal side length ratios. That being said, the point $E$ gives purposes to the midpoint $L$ of $AB$. To complete the proof, we need to prove $\angle AIK=\angle BYP$ and $\angle AIL=\angle CPX$. To this end, we need to prove that triangle $AKI$ is similar to triangle $BPY$ and triangle $ALI$ is similar to triangle $CPX$, which is done by side length ratio chasing, which is obtained from the pairs of similar triangles above. A full solution is published at https://storage.googleapis.com/deepmind-media/DeepMind.com/Blog/imo-2024-solutions/P4/index.html. This solution was obtained within 30 seconds at IMO 2024 and was given the full seven points by Joseph Myers, a two-time IMO gold medalist and Chair of the IMO 2024 Problem Selection Committee.
@@
Along with IMO 2024 P4, AlphaGeometry can solve many challenging problems with only 1 extra auxiliary point, some of which involve rather unconventional constructions. One such problem is IMO 2013 P3.
@@
IMO 2014 P3: Convex quadrilateral ABCDABCD has $\angle ABC=\angle CDA=90^{\circ}$. Point $H$ is the foot of the perpendicular from $A$ to $BD$. Points $S$ and $T$ lie on sides ABAB and $AD$, respectively, such that $H$ lies inside triangle $SCT$ and $\angle CHS-\angle CSB=90^{\circ},\angle THC-\angle DTC=90^{\circ}.$ Prove that line $BD$ is tangent to the circumcircle of triangle $TSH$.
@@
Our final example is the G7 problem from the IMO 2009 Shortlist.
@@
To the best of our knowledge, this problem previously had only computational solutions, e.g., by using complex numbers, trigonometric computations or proof by contradiction via an inequality argument. Since AlphaGeometry does not have access to these computational and reasoning tools, as well as advanced Euclidean geometry knowledge, we originally expected that this problem cannot be solved by AlphaGeometry. Nevertheless, AlphaGeometry was able to produce an elegant solution with only angle and ratio chasing by constructing key auxiliary constructions. First, AlphaGeometry shows that $X$ and $Z$ are reflections of each other w.r.t. $BI$, and by symmetry it follows that $I$ is the circumcenter of $\triangle XYZ$. From this we can show that $AB=AC$, and by symmetry, we have $\triangle ABC$ is an equilateral triangle. However, the main challenge with this problem is to use the condition $\triangle XYZ$ being an equilateral triangle, i.e., $XY=YZ$ and its cyclic variants. To this end, AlphaGeometry constructs a series of circumcenters of key triangles:
@@
Appendix E Additional evaluation on the hardest IMO shortlist problems
@@
Appendix F Towards generating full proofs with a language model
@@
  • Unverified: The premises of the step do not imply the conclusion of the step under the deduction rule that it uses.
@@
For evaluation, we query our language models with the 2000-2024 IMO problems and 32 samples at temperature 1.0. The models are queried without any end-of-sentence tokens such that they generate full proofs. Then we compute how many valid proof steps the models produce on average across samples and all problems. It turns out our models do not make many syntax mistakes (see Figure 16), the majority of generated steps are either valid (either fully verified or correct but unverified). One surprising find is that both small and larger models perform similarly. These results support the idea that large language models can be self-sufficient without depending on external tools, but until inference speed is improved and hallucinations are completely resolved, the tools will stay essential for math applications.
@@
Appendix G Automated problem formalization and diagram generation
In this section we describe our progress in building a fully-automated AG2 system taking input in natural language and outputting the full proof along with automatically constructed diagrams. This system was used at IMO 2025 to solve the geometry problem (P2) from natural language in 20 seconds.
Automated formalization.
A major weakness of AlphaGeometry, and other similar neuro-symbolic systems, is the need to manually transform input problems from natural language into a domain-specific language. For example, a simple geometry problem in natural language, “Given a triangle $ABC$ with two equal sides $AB=AC$, prove that angles $B$ and $C$ are equal”, becomes triangle a b c; a b = a c ? eqangle b a b c c b c a in the AlphaGeometry domain language.
Automating this process, called formalization, is an active area of research (see for example, szegedy2020promising; wu2022autoformalizationlargelanguagemodels; jiang2023multilingual; poiroux2024improving; murphykaiyu2024autoformalizinggeometry). It is a significantly more complicated problem compared to a translation between human languages. While translation aims to preserve meaning, formalization frequently requires re-formulating the original problem into an alternative form, and sometimes disambiguating the nuances in the original problem statement. Automated formalization (auto-formalization), therefore, demands significant background knowledge and problem-solving skills on its own. Given that recently foundation models started demonstrating such capabilities, we use one such model, Gemini team2024gemini, to automate the problem formalization for AlphaGeometry. We start by manually translating several dozens of geometry problems into the AG language. Then we use these examples to write a few-shot prompt asking Gemini to translate a given geometry problem from natural language into the AG language. We query Gemini five times with this prompt, followed by another Gemini call asking to combine these results into one final answer. With this approach, we are able to formalize 33 out of 44 formalizable IMO 2000-2024 geometry problems. For easier geometry problems, it is very consistent and makes almost no mistakes.
Automated diagram generation.
Another manual part of our pipeline was diagram generation. In AG1, each point is defined by at most two basic predicates recalled in Table 1, the problem is therefore defined constructively and diagrams can be generated automatically. In AG2, we allow one or multiple points to be defined simultaneously by an arbitrary number of predicates, allowing us to also cover non-constructive problems. Consider a non-constructive problem statement, “Let $ABC$ be a triangle with incenter $I$, such that $IA=2IB$ …”, here point $I$ is not only defined as an incenter, i.e. the intersection of two internal bisectors, but also defined by a third predicate $IA=2IB$ and there is no general strategy to construct such four points. Since AG2 covers non-constructive problems, diagram construction becomes a non-trivial part of the pipeline and generally requires human intervention. Similar to krueger2021automatically, below we propose a new algorithm to automatically generate diagrams given non-constructive problem specifications.
We start by initializing the points. We have three initialization methods, which we alternate in different attempts:
  1. 1.
  2. 2.
    Construction in order.
  3. 3.
    Construction in a heuristically chosen order.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
This three-stage optimization method is similar to the methodology introduced in krueger2021automatically. The final stage addresses the practical limitations encountered when tuning the gradient descent optimization in the original method, where achieving a consistently satisfactory error margin proved challenging.
Appendix H Inequality rules
This section details a set of inequality rules. Although potentially useful for many inequality-related questions from Euclidean geometry, these rules were not included in the AlphaGeometry2 system.
H.1 Definitions
  • $P(A_{1}A_{2}\ldots A_{n})$ is a polygon formed by $A_{1},A_{2},\ldots,A_{n}$.
  • $\omega(ABC)=1$ if $ABC$ is clockwise, 1-1 otherwise.
  • $\eta(ABC)=1$ if $ABC$ collinear and $B$ is between $A$ and $C$, 1-1 if $A$ and $C$ are on the same side with respect to $B$.
  • $\angle ABC$ the angle formed by the segments $AB$ and $BC$ whose measure is less than $\pi$.
  • $\angle_{0}(ABC)$: traditional unoriented measure of $\angle ABC$ Ranges from 0 to $\pi$.
  • $\angle_{1}(ABC)$: angle measured from line $AB$ to $BC$ counterclockwise. Ranges from 0 to $\pi$.
  • $\angle_{2}(ABC)$: angle measured from ray $BA$ to $BC$. Positive if counterclockwise and negative if clockwise. Ranges from $\pi$ to $\pi$.
H.2 How to write some of the common geometric inequality statements using $\omega$
  • $C$ and $D$ separated by $AB$: $\omega(ABC)=\omega(ADB)$
  • $D$ is contained in $P(ABC)$: $\omega(DBC)=\omega(ADC)=\omega(ABD)(=\omega(ABC))$
  • $P(A_{1}A_{2}\ldots A_{n})$ is convex: $\omega(A_{1}A_{2}A_{3})=\omega(A_{2}A_{3}A_{4})=\ldots=\omega(A_{n}A_{1}A_{2})$.
  • $p\in P(A_{1}A_{2}\ldots A_{n})$ is convex: $\omega(pA_{1}A_{2})=\omega(pA_{2}A_{3})=\ldots=\omega(pA_{n}A_{1})(=\omega(A_{1}A_{2}A_{3})=\omega(A_{2}A_{3}A_{4})=\ldots=\omega(A_{n}A_{1}A_{2}))$
  • $D$ is in the (minor) sector determined by the ray $AB$ and ray $AC$: $\omega(BAD)=\omega(BAC)=\omega(DAC)$
H.3 Trivial rules
  • ncol $A\penalty 10000\ B\penalty 10000\ C$ $\Rightarrow$ $\omega(ABC)=\omega(BCA)$
  • ncol $A\penalty 10000\ B\penalty 10000\ C$ $\Rightarrow$ $\omega(ABC)\neq\omega(ACB)$
  • $\Rightarrow$ $\eta(ABC)=\eta(CBA)$
  • $\omega(ABC)\neq\omega(DEF)$ $\Rightarrow$ $\omega(ABC)=\omega(DFE)$
  • $\Rightarrow$ $\angle_{0}(ABC)=\omega(ABC)\angle_{2}(ABC)$
  • $\Rightarrow$ $\angle_{0}(ABC)=|\angle_{2}(ABC)|$
  • $\Rightarrow$ $\angle_{1}(ABC)=\angle_{2}(ABC)+\pi(1-\omega(ABC))/2$
H.4 Rules for relating $\omega$ and $\eta$
  • $\eta(ABC)=1$, ncol $X\penalty 10000\ A\penalty 10000\ B$ $\Rightarrow$ $\omega(XAB)=\omega(XBC)=\omega(XAC)$
  • $\eta(ABC)=-1$, ncol $X\penalty 10000\ A\penalty 10000\ B$ $\Rightarrow$ $\omega(XBA)=\omega(XBC)$
  • ncol $X\penalty 10000\ A\penalty 10000\ B$, col $A\penalty 10000\ B\penalty 10000\ C$, $\omega(XAB)=\omega(XBC)$ $\Rightarrow$ $\eta(ABC)=1$
  • ncol $X\penalty 10000\ A\penalty 10000\ B$, col $A\penalty 10000\ B\penalty 10000\ C$, $\omega(XBA)=\omega(XBC)$ $\Rightarrow$ $\eta(ABC)=-1$
H.5 Basic rules using P in polygon
  • ncol $A_{1}\penalty 10000\ A_{2}\penalty 10000\ A_{3}$ $\Rightarrow$ $P(A_{1}A_{2}A_{3})$: convex
  • $\omega(A_{1}A_{2}A_{3})=\omega(A_{2}A_{3}A_{4})=\ldots=\omega(A_{n}A_{1}A_{2})$ $\Rightarrow$ $P(A_{1}A_{2}\ldots A_{n})$: convex
  • $P(A_{1}A_{2}\ldots A_{n})$: convex $\Rightarrow$ $\omega(A_{1}A_{2}A_{3})=\omega(A_{i}A_{j}A_{k})$ with any $i<j<k$
  • $P(A_{1}A_{2}\ldots A_{n})$: convex, $\omega(pA_{1}A_{2})=\omega(pA_{2}A_{3})=\ldots=\omega(pA_{n}A_{1})$ $\Rightarrow$ $p\in P(A_{1}A_{2}\ldots A_{n})$
  • $p\in P(A_{1}A_{2}\ldots A_{n})$ $\Rightarrow$ $\omega(pA_{1}A_{2})=\omega(A_{1}A_{2}A_{3})$
  • $p\in P(A_{1}A_{2}\ldots A_{n})$, $P(A_{1}A_{2}\ldots A_{n}A_{n+1})$: convex $\Rightarrow$ $p\in P(A_{1}A_{2}\ldots A_{n}A_{n+1})$
  • $p\in P(ABC)$, $\eta(DBC)=1$ $\Rightarrow$ $p\in P(ADC)$
  • $p\in P(A_{1}A_{2}\ldots A_{n})$, col $X\penalty 10000\ A_{n}\penalty 10000\ A_{1}$, col $X\penalty 10000\ A_{n-1}\penalty 10000\ A_{n-2}$ $\Rightarrow$ $p\in P(A_{1}A_{2}\ldots A_{n-2}X)$
  • $P(A_{1}\ldots A_{n})$: convex $\Rightarrow$ $P(A_{1}\ldots A_{n}-1)$: convex
  • $P(ABCD)$: convex, col $A\penalty 10000\ X\penalty 10000\ C$, col $B\penalty 10000\ X\penalty 10000\ D$ $\Rightarrow$ $X\in P(ABCD)$
H.6 Basic rules using $\in O$” and $\notin O$
Here, $O(P,AB)$ is a circle centered at $P$ with the radius $AB$, and $O(ABC)$ is the circumcircle of $P(ABC)$.
  • col $A\penalty 10000\ O\penalty 10000\ B$, $OA=OB$, $\eta(AOB)=1$, $\angle_{0}ACB<\pi/2$ $\Rightarrow$ $C\notin O(O,AO)$
  • col $A\penalty 10000\ O\penalty 10000\ B$, $OA=OB$, $\eta(AOB)=1$, $\angle_{0}ACB>\pi/2$ $\Rightarrow$ $C\in O(O,AO)$
  • $\omega(ADB)=\omega(ACB)$, $\angle_{0}ACB>\angle_{0}ADB$ $\Rightarrow$ $D\notin O(ABC)$
  • $\omega(ADB)=\omega(ACB)$, $\angle_{0}ACB<\angle_{0}ADB$ $\Rightarrow$ $D\in O(ABC)$
  • $\omega(ADB)=-\omega(ACB)$, $\pi-\angle_{0}ACB>\angle_{0}ADB$ $\Rightarrow$ $D\notin O(ABC)$
  • $\omega(ADB)=-\omega(ACB)$, $\pi-\angle_{0}ACB<\angle_{0}ADB$ $\Rightarrow$ $D\in O(ABC)$
  • $D\notin O(ABC)$, $\omega(ADB)=\omega(ACB)$ $\Rightarrow$ $\angle_{0}ACB>\angle_{0}ADB$
  • $D\notin O(ABC)$, $\omega(ADB)=-\omega(ACB)$ $\Rightarrow$ $\pi-\angle_{0}ACB>\angle_{0}ADB$
  • $p\in P(ABC)$ $\Rightarrow$ $p\in O(ABC)$
  • $p\in O(ABC)$, $\omega(pBC)=\omega(ABC)$ $\Rightarrow$ $\angle_{0}BpC>\angle_{0}BAC$
  • $p\in O(ABC)$, $\omega(pBC)=-\omega(ABC)$ $\Rightarrow$ $\angle_{0}BpC>\pi-\angle_{0}BAC$
  • $D\in O(O,OA)$ $\Rightarrow$ $OD<OA$
  • $OD<OA$ $\Rightarrow$ $D\in O(O,OA)$
  • $D\notin O(O,OA)$ $\Rightarrow$ $OD>OA$
  • $OD>OA$ $\Rightarrow$ $D\notin O(O,OA)$
H.7 Acute and obtuse angles
  • $\angle ABC$: acute $\Rightarrow$ $\angle_{0}ABC<\pi/2$
  • $\pi>\angle_{2}(ABC)>\pi/2$, circ $O\penalty 10000\ A\penalty 10000\ B\penalty 10000\ C$ $\Rightarrow$ $\omega(OAC)=\omega(BCA)$
  • $\pi>\angle_{2}(ABC)>\pi/2$, $H$: orthocenter of $P(ABC)$ $\Rightarrow$ $\omega(AHC)=\omega(ABC)$
  • $G$: centroid of $P(ABC)$ $\Rightarrow$ $G\in P(ABC)$
  • II: incenter of $P(ABC)$ $\Rightarrow$ $I\in P(ABC)$
  • $P(ABC)$: acute $\Rightarrow$ $O,H\in P(ABC)$
  • col $B\penalty 10000\ D\penalty 10000\ C$, ncol $A\penalty 10000\ B\penalty 10000\ D$, $AB=AD$, $\angle ABC$: acute $\Rightarrow$ $\eta(DBC)=-1$
  • col $B\penalty 10000\ D\penalty 10000\ C$, ncol $A\penalty 10000\ B\penalty 10000\ D$, $AB=AD$, $\angle ABC$: obtuse $\Rightarrow$ $\eta(DBC)=1$
  • col $B\penalty 10000\ D\penalty 10000\ C$, $\angle_{0}(ABC)<\angle_{0}(ADC)$ $\Rightarrow$ $\eta(DBC)=-1$
  • midpoint $M\penalty 10000\ B\penalty 10000\ C$, col $A\penalty 10000\ X\penalty 10000\ B$, $XM\perp BC$, $AB>AC$ $\Rightarrow$ $\eta(AXB)=1$
  • midpoint $M\penalty 10000\ B\penalty 10000\ C$, col $A\penalty 10000\ X\penalty 10000\ B$, $XM\perp BC$, $AB<AC$ $\Rightarrow$ $\eta(XAB)=1$
H.8 Some inequalities
  • $\eta(ABC)=1$ $\Rightarrow$ $AC>AB$
  • $\eta(ABC)=1$ $\Rightarrow$ $AC=AB+BC$
  • col $A\penalty 10000\ B\penalty 10000\ C$, $AC>AB$, $AC>BC$ $\Rightarrow$ $\eta(ABC)=1$
  • $\omega(BAD)=\omega(BAC)=\omega(DAC)$ $\Rightarrow$ $\angle_{0}(BAD)<\angle_{0}(BAC)$, $\angle_{0}(DAC)<\angle_{0}(BAC)$
  • $\angle_{0}(BAD)<\angle_{0}(BAC)$, $\angle_{0}(DAC)<\angle_{0}(BAC)$ $\Rightarrow$ $\omega(BAD)=\omega(BAC)=\omega(DAC)$
  • $\Rightarrow$ $AB+BC\geq AC$.
  • $AB+BC=AC$ $\Rightarrow$ col $A\penalty 10000\ B\penalty 10000\ C$, $\eta(ABC)=1$
  • col $B\penalty 10000\ C\penalty 10000\ D$, col $A\penalty 10000\ O\penalty 10000\ D$, $OA=OB$, $OB=OC$, $P(ABC):$ acute, $AB<AC$ $\Rightarrow$ $BD>DC$
  • col $B\penalty 10000\ C\penalty 10000\ D$, col $A\penalty 10000\ O\penalty 10000\ D$, $OA=OB$, $OB=OC$, $\angle_{0}(BAC)>\pi/2$, $AB<AC$ $\Rightarrow$ $BD<DC$
  • col $A\penalty 10000\ B\penalty 10000\ C$, $\eta(BAC)=-1$, $CA>BA$ $\Rightarrow$ $\eta(ABC)=1$
  • $\eta(ACB)=1$, $\eta(ADB)=1$ $\Rightarrow$ $\eta(CAD)=-1$
  • $AB\perp BC$ $\Rightarrow$ $AC>AB$, $AC>BC$
  • $AB>AC$ $\Rightarrow$ $\angle_{0}(ACB)>\angle_{0}(ABC)$
  • $\angle_{0}(ACB)>\angle_{0}(ABC)$ $\Rightarrow$ $AB>AC$
H.9 General $\omega$ and $\eta$ deduction rules
  • $\omega(ABC)=\omega(BCD)$, $\omega(BCD)=\omega(CDA)$ $\Rightarrow$ $\omega(DAB)=\omega(CDA)$
  • $\angle_{2}(ABC)>0$ $\iff$ $\omega(ABC)=1$
  • $\angle_{2}(ABC)<0$ $\iff$ $\omega(ABC)=-1$
  • $\Rightarrow$ $\angle_{0}(ABC)+\angle_{0}(BCA)+\angle_{0}(CAB)=\pi$
  • $\eta(BCX)=1$ $\Rightarrow$ $\angle_{2}(ACX)=\angle_{2}(ABC)+\angle_{2}(CAB)$
  • circ $O\penalty 10000\ A\penalty 10000\ B\penalty 10000\ C$, $\omega(BOA)=\omega(COB)$ $\Rightarrow$ $\omega(ABC)=\omega(BOA)$
  • $\eta(CAB)=-1$, $\eta(ACB)=-1$ $\Rightarrow$ $\eta(ABC)=1$
  • $\eta(ABC)=1$ $\Rightarrow$ $\eta(CAB)=-1$, $\eta(ACB)=-1$
  • $\eta(ABC)=1$, $\eta(BCD)=1$ $\Rightarrow$ $\eta(ABD)=1$, $\eta(ACD)=1$
  • col $B\penalty 10000\ X\penalty 10000\ C$, $AX\perp BC$, $\angle_{0}(ABC)<\pi/2$, $\angle_{0}(ACB)<\pi/2$ $\Rightarrow$ $\eta(BXC)=1$
  • col $B\penalty 10000\ X\penalty 10000\ C$, $\angle_{2}(BAX)=\angle_{2}(XAC)$ $\Rightarrow$ $\eta(BXC)=1$
  • midpoint $M\penalty 10000\ A\penalty 10000\ B$ $\Rightarrow$ $\eta(AMB)=1$
  • $\eta(ADB)=\eta(AEC)=1$ $\Rightarrow$ $P(DECB)$: convex
  • $X\notin O(ABC)$, $\omega(XAB)=\omega(CAB)$, cyclic $A\penalty 10000\ B\penalty 10000\ C\penalty 10000\ D$, col $X\penalty 10000\ A\penalty 10000\ D$, $\angle_{0}(CBA)>\angle_{0}(CAX)$ $\Rightarrow$ $\eta(XDA)=1$
  • $X\notin O(ABC)$, $\omega(XAB)=\omega(CAB)$, cyclic $A\penalty 10000\ B\penalty 10000\ C\penalty 10000\ D$, col $X\penalty 10000\ A\penalty 10000\ D$, $\angle_{0}(BCA)+\angle_{0}(XAB)<\pi$ $\Rightarrow$ $\eta(XDA)=1$
  • $X\notin O(ABC)$, $\omega(XAB)=\omega(CAB)$, cyclic $A\penalty 10000\ B\penalty 10000\ C\penalty 10000\ D$, col $X\penalty 10000\ A\penalty 10000\ D$, $\angle_{0}(CBA)<\angle_{0}(CAX)$ $\Rightarrow$ $\eta(XAD)=1$
  • $X\notin O(ABC)$, $\omega(XAB)=\omega(CAB)$, cyclic $A\penalty 10000\ B\penalty 10000\ C\penalty 10000\ D$, col $X\penalty 10000\ A\penalty 10000\ D$, $\angle_{0}(BCA)+\angle_{0}(XAB)>\pi$ $\Rightarrow$ $\eta(XAD)=1$
  • $X\notin O(ABD)$, col $A\penalty 10000\ O\penalty 10000\ B$, col $X\penalty 10000\ A\penalty 10000\ D$, $\angle_{0}(XAB)<\pi/2$ $\Rightarrow$ $\eta(XDA)=1$
  • $X\notin O(ABD)$, col $A\penalty 10000\ O\penalty 10000\ B$, col $X\penalty 10000\ A\penalty 10000\ D$, $\angle_{0}(XAB)>\pi/2$ $\Rightarrow$ $\eta(XAD)=1$
  • $X\in O(ABC)$, col $A\penalty 10000\ X\penalty 10000\ B$ $\Rightarrow$ $\eta(AXB)=1$
  • $\angle_{2}XBA+\angle_{2}XBC=\omega(XBA)\pi$, $\angle_{2}XCA+\angle_{2}XCB=\omega(XCA)\pi$ $\Rightarrow$ $X\notin O(ABC)$ ($X$ will be the excenter relative to $A$: $X==I_{A}$)